OpenStudy (anonymous):

A playground is on the flat roof of a city school, 6.00 m above the street below. The vertical wall of the building is 7.00 m high, forming a 1.00 m high railing around the playground. A ball has fallen to the street below, and a passerby returns it by launching it at an angle of 53.0° above the horizontal at a point 24.0 m from the base of the building wall. The ball takes 2.20 s to reach a point vertically above the wall. (a) Find the speed at which the ball was launched. (b) Find the vertical distance by which the ball clears the wall. (c) Find the horizontal distance from the wall to the p

OpenStudy (lastdaywork):



OpenStudy (lastdaywork):

The diagram somehow got shifted towards right..the part that got missing is only labeling of heights. Now, can you post your attempt for the question ??

OpenStudy (anonymous):

where is the answer?

OpenStudy (lastdaywork):

Why don't you show me your attempt first ?? Whatever equations you made..or anything..

OpenStudy (anonymous):

here is the diagram

OpenStudy (lastdaywork):

Look, if you are just looking for an answer to your question; then I can't help you...its also against the Code of Conduct of this website. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i get first the initial velocity

OpenStudy (anonymous):

Vi=24/2.2cos53

OpenStudy (anonymous):

Vi=18.1m/s

OpenStudy (lastdaywork):

" (c) Find the horizontal distance from the wall to the p" Can you complete the question ??

OpenStudy (anonymous):

after that i don't know what to do

OpenStudy (anonymous):

. (c) Find the horizontal distance from the wall to the point on the roof where the ball lands.

OpenStudy (lastdaywork):

Okay.. The calculation you did for finding velocity is wrong..d < range of the projectile. Hence you CANNOT write : Vcos(53)*2.2 = 24

OpenStudy (anonymous):

what is the correct?

OpenStudy (anonymous):

just give me the formulas

OpenStudy (lastdaywork):

Can you write the equations for the motion in x and y direction ??

OpenStudy (anonymous):

i don't know that

OpenStudy (anonymous):

just give to me the formulas

OpenStudy (lastdaywork):

The usual equations are - y = Vsin(53)t - (1/2)gt^2 x = Vcos(53)t Vy = Vsin(53) - gt Vx = Vcos(53) Find the time when y = 7 ; you'll get two values of time ; let the bigger one be T Then, Vcos(53)T = 24 ; you can find V from here Next; find time when x = 24; substitute it in y = Vsin(53)t - (1/2)gt^2 to get ; then (y - 7) is the answer for part (b) Answer for part (c) is ( Vcos(53)*2.2 - 24)

OpenStudy (anonymous):

what is V stands for?

OpenStudy (lastdaywork):

V is the magnitude of initial velocity vector.